# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ 3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
We use the standard “continuity” analysis of Cohen-type symmetric models. In the basic Cohen model $V(A)$ generated by an unordered set $A$ of Cohen reals, a key lemma (sometimes called a continuity or indiscernibility lemma) shows that statements over $V(A)$ with parameters from $V$ are already decided in a subextension generated by finitely many members of $A$; see, e.g., Felgner [4, p. 133] or Blass [3, Prop. 1.2] (and the discussion in [6, Lem. 2.8]).
More precisely, for each $x\in X$ there is a $\mathbb{P}$–name $\tau_{x}\in V$ and a sequence $s_{x}\in A^{\omega}$ (coding the countable symmetry support relevant to $\tau_{x}$) such that $x$ is recovered uniformly from the pair $(\tau_{x},s_{x})$. Enumerating the set of $\mathbb{P}$–names in $V$ by ordinals yields the required surjection from $A^{\omega}\times\eta$ onto $X$ for some ordinal $\eta$.
The same argument applies to our Cohen symmetric model $\mathcal{N}$ built from $\operatorname{Add}(\omega,\omega_{1})$ with the countable-support symmetry filter: every set $X\in\mathcal{N}$ is obtained (uniformly in $X$) from $V$ and a countable sequence of members of $A$, and therefore is a surjective image of $A^{\omega}\times\eta$ for some ordinal $\eta$. Equivalently, $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$. ∎
@@ 4 Reduction blueprint: global $\mathsf{PP}$ from localized $\mathsf{PP}$
For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [17, §2.1, Fact]
@@
Assume $\mathsf{SVC}^{+}(T)$. Then
@@
[17, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$). As this argument is critical to our goals, we demonstrate a proof here.
@@
($\Rightarrow$) If $\mathsf{PP}$ holds, then $\mathsf{PP}\!\restriction T$ is immediate. Moreover, $\mathsf{PP}$ implies $\mathsf{AC}_{\mathsf{WO}}$ (e.g. via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [17, Proposition 3.17 and footnote 4] for references).
($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\pi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
@@
$\pi_{X}:X\to T\times\eta,\quad\pi_{X}(x)=j(\langle x,0\rangle),\qquad\pi_{Y}:Y\to T\times\eta,\quad\pi_{Y}(y)=j(\langle y,1\rangle),$
@@
$A:=\operatorname{ran}(\pi_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\pi_{Y})\subseteq T\times\eta.$
Then $\pi_{X}:X\to A$ and $\pi_{Y}:Y\to B$ are bijections, hence have inverses $\pi_{X}^{-1}:A\to X$ and $\pi_{Y}^{-1}:B\to Y$. Define
$f\;:=\;\pi_{X}\circ\pi\circ\pi_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
@@
$i\;:=\;\pi_{Y}^{-1}\circ g\circ\pi_{X}\;:\;X\to Y$
@@
Proposition 4.7 (The $\mathsf{AC}_{\mathsf{WO}}$ conjunct cannot be omitted).
There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
$M\models\mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
In particular, global $\mathsf{PP}$ cannot be recovered from a small-choice hypothesis together with a single-parameter instance $\mathsf{PP}(S^{\ast})$.
Reference.
This is witnessed by Cohen’s first model; see [17, Proposition 3.22]. ∎
Proposition 4.8 (The localized $\mathsf{PP}\!\restriction S$ component cannot be omitted).
There is a model $M\models\mathsf{ZF}$ and a set $S^{\ast}\in M$ such that
$M\models\mathsf{AC}_{\mathsf{WO}}\ \wedge\ \mathsf{SVC}^{+}(S^{\ast})\ \wedge\ \mathsf{PP}(S^{\ast})\ \wedge\ \neg\mathsf{PP}.$
Consequently, even assuming $\mathsf{AC}_{\mathsf{WO}}$ and small-choice hypotheses, one cannot in general replace the localized requirement $\mathsf{PP}\!\restriction S^{\ast}$ by the weaker single-parameter assertion $\mathsf{PP}(S^{\ast})$ and still obtain global $\mathsf{PP}$.
Reference.
This is witnessed by the Feferman-style model $N_{\aleph_{1}}$ of Truss; see [17, Proposition 3.23]. ∎
Remark 4.9 (Why we keep both conjuncts).
The preceding propositions are included only to justify (conceptually) why the localization theorem is stated with both components present: the “choice” side $\mathsf{AC}_{\mathsf{WO}}$ and the “localized comparability” side $\mathsf{PP}\!\restriction T$. We will force both conjuncts $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ in Theorem 4.5. In Ryan–Smith’s analysis the corresponding reduction is sharp: in general one cannot drop the displayed conjuncts (nor substantially weaken the hypotheses) without losing the equivalence; see [17, Props. 3.22 and 3.23].
@@ 4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (in practice, $M$ will be an intermediate stage of our symmetric iteration). Work with the fixed parameter $T=\mathcal{P}(S)$ from Definition 3.27.
@@
(see [17, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.71.
(We also note that Ryan–Smith’s localization for $\mathsf{PP}$ uses both conjuncts: $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$; see [17, Prop. 3.17] and Question 3.18 therein, and see [17, Props. 3.22 and 3.23] for counterexamples showing the conjuncts cannot in general be omitted.)
@@
$\mathbb{Q}_{f}$ is Countably-closed (sometimes called $\omega_{1}$-closed).
@@
The $\omega_{1}$-closure of $\mathbb{Q}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”. In a bare $\mathsf{ZF}$ context this need not hold, so the closure claim should be read as relative to our standing hypothesis $M\models\mathsf{ZF}+\mathsf{DC}$.
@@
Let $M\models\mathsf{ZF}+\mathsf{DC}$ be transitive and let $\mathbb{P}\in M$ be countably closed in $M$ (i.e. every decreasing $\omega$-sequence in $M$ has a lower bound in $M$). Then:
@@
    $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
@@
    Forcing with $\mathbb{P}$ over $M$ preserves $\mathsf{DC}$.
@@
(1) No new $\omega$-sequences of ordinals. Let $p_{0}\in\mathbb{P}$ and $\dot{a}\in M$ with $p_{0}\Vdash\dot{a}:\check{\omega}\to\mathrm{Ord}$. For each $n<\omega$, let
$D_{n}:=\Bigl\{p\leq p_{0}:\exists\alpha\in\mathrm{Ord}\ (p\Vdash\dot{a}(\check{n})=\check{\alpha})\Bigr\}.$
Each $D_{n}$ is dense below $p_{0}$.
@@
$\Sigma:=\Bigl\{(n,p,\vec{\alpha}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \vec{\alpha}\in\mathrm{Ord}^{n},\ p\Vdash\forall k<n\ \dot{a}(\check{k})=\check{\vec{\alpha}(k)}\Bigr\}.$
@@
$(n,p,\vec{\alpha})\ R\ (n+1,q,\vec{\beta})$
iff $q\leq p$, $\vec{\beta}\in\mathrm{Ord}^{n+1}$ extends $\vec{\alpha}$, and $q\Vdash\dot{a}(\check{n})=\check{\vec{\beta}(n)}$. Then $R$ is total on $\Sigma$ by density of $D_{n}$.
By $\mathsf{DC}$ in $M$, there is an $R$-chain $\langle(n,p_{n},\vec{\alpha}_{n}):n<\omega\rangle\in M$ starting from some $(0,p_{0},\emptyset)\in\Sigma$ (note $\mathrm{Ord}^{0}=\{\emptyset\}$). By countable closure, there is $p_{\omega}\in\mathbb{P}$ with $p_{\omega}\leq p_{n}$ for all $n$.
Let $g:=\bigcup_{n<\omega}\vec{\alpha}_{n}$. Since the chain is an element of $M$, Replacement in $M$ yields $g\in M$ and $g:\omega\to\mathrm{Ord}$. By construction, $p_{\omega}\Vdash\forall n\ \dot{a}(\check{n})=\check{g(n)}$, hence $p_{\omega}\Vdash\dot{a}=\check{g}$. Thus $\mathbb{P}$ adds no new $\omega$-sequences of ordinals.
(2) Preservation of $\mathsf{DC}$. Let $p_{0}\in\mathbb{P}$ and $\dot{X},\dot{R},\dot{x}_{0}\in M$ with
$p_{0}\Vdash\ \dot{x}_{0}\in\dot{X}\ \wedge\ (\forall x\in\dot{X})(\exists y\in\dot{X})\ x\,\dot{R}\,y.$
@@
$\Gamma:=\Bigl\{(n,p,\dot{x}):n<\omega,\ p\in\mathbb{P},\ p\leq p_{0},\ \dot{x}\in M,\ p\Vdash\dot{x}\in\dot{X}\Bigr\}.$
@@
$(n,p,\dot{x})\ S\ (n+1,q,\dot{y})$
iff $q\leq p$ and $q\Vdash\dot{y}\in\dot{X}\ \wedge\ \dot{x}\,\dot{R}\,\dot{y}$. Totality of $S$ follows from the premise forced by $p_{0}$.
By $\mathsf{DC}$ in $M$, pick an $S$-chain $\langle(n,p_{n},\dot{x}_{n}):n<\omega\rangle\in M$ starting with $(0,p_{0},\dot{x}_{0})\in\Gamma$. By countable closure, let $p_{\omega}\leq p_{n}$ for all $n$.
@@
$\dot{f}:=\bigl\{\langle\langle\check{n},\dot{x}_{n}\rangle,\mathbbm{1}_{\mathbb{P}}\rangle:n<\omega\bigr\}.$
Since the chain is in $M$, Replacement gives $\dot{f}\in M$. Then $p_{\omega}\Vdash\dot{f}:\check{\omega}\to\dot{X}$ and $p_{\omega}\Vdash\forall n\ \dot{f}(\check{n})\,\dot{R}\,\dot{f}(\check{n+1})$, which is exactly $\mathsf{DC}$ in the extension. ∎
@@
Apply Lemma 4.18 with $\mathbb{P}=\mathbb{Q}_{f}$. Countable closure holds by Lemma 4.15. ∎

### v4

@@ 3.5 A safe $\mathsf{SVC}$ seed in $\mathcal{N}$
Definition 3.25 ($\mathsf{SVC}(S)$).
Let $S$ be a set. We write $\mathsf{SVC}(S)$ for the statement:
$\forall X\,\exists\eta\in\mathrm{Ord}\,\exists e\ (e:S\times\eta\twoheadrightarrow X).$
@@
We work in $\mathcal{N}$. The conclusion $\mathsf{SVC}(A^{\omega})$ is a standard consequence of the “continuous reading/continuity” analysis for generalized Cohen symmetric models under a symmetry filter generated by $\operatorname{Fix}(E)$ for countable $E\subseteq\omega_{1}$.
In the finite-support Cohen model $V(A)$, the corresponding analysis yields $\mathsf{SVC}(A^{<\omega})$ (see Felgner [4, p. 133] and Blass [3, Prop. 1.2]). In our present seed model $\mathcal{N}$, the symmetry filter is generated by $\operatorname{Fix}(E)$ for countable $E$, and thus the same argument runs with $\omega$-sequences of elements of $A$ in place of finite tuples. Concretely, the relevant input is exactly:
  1. 1.
    every hereditarily symmetric name admits a countable symmetry support (Lemma 3.14); and
  2. 2.
    every countable $E\subseteq\omega_{1}$ from $V$ can be coded by an element of $A^{\omega}$ in $\mathcal{N}$ (Lemma 3.23).
A detailed presentation of this countable-support version is given in [6, Lem. 2.8]. Therefore $\mathcal{N}\models\mathsf{SVC}(A^{\omega})$. ∎
@@ 4 Reduction blueprint: global $\mathsf{PP}$ from localized $\mathsf{PP}$
The notation $\mathsf{PP}\!\restriction T$ follows [18, §3.8] and is unrelated to forcing-condition restrictions $p\restriction E$.
@@
For every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$. [18, §2.1, Fact]
@@
Assume $\mathsf{SVC}^{+}(T)$ (Definition 4.2). Then
@@
[18, Proposition 3.17] (stated there with parameter “SS”; we apply it with $S:=T$). Here $\mathsf{PP}\!\restriction T$ is as in Definition 4.1, and $\mathsf{AC}_{\mathsf{WO}}$ is as in Remark $1.$ As this argument is critical to our goals, we demonstrate a proof here.
@@
($\Rightarrow$) If $\mathsf{PP}$ holds, then $\mathsf{PP}\!\restriction T$ is immediate. Moreover, $\mathsf{PP}$ implies $\mathsf{AC}_{\mathsf{WO}}$ (e.g. via the consequence $(\forall X)\,\aleph(X)=\aleph^{\ast}(X)$; see [18, Proposition 3.17 and footnote 4] for references).
($\Leftarrow$) Assume $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, and let $\varphi:Y\twoheadrightarrow X$ be a surjection. We must produce an injection $X\hookrightarrow Y$.
@@
$\iota_{X}:X\to T\times\eta,\quad\iota_{X}(x)=j(\langle x,0\rangle),\qquad\iota_{Y}:Y\to T\times\eta,\quad\iota_{Y}(y)=j(\langle y,1\rangle),$
@@
$A:=\operatorname{ran}(\iota_{X})\subseteq T\times\eta,\qquad B:=\operatorname{ran}(\iota_{Y})\subseteq T\times\eta.$
Then $\iota_{X}:X\to A$ and $\iota_{Y}:Y\to B$ are bijections, hence have inverses $\iota_{X}^{-1}:A\to X$ and $\iota_{Y}^{-1}:B\to Y$. Define
$f\;:=\;\iota_{X}\circ\varphi\circ\iota_{Y}^{-1}\;:\;B\twoheadrightarrow A.$
@@
$i\;:=\;\iota_{Y}^{-1}\circ g\circ\iota_{X}\;:\;X\to Y$
@@
Remark 4.7 (On the hypotheses in Theorem 4.5).
Theorem 4.5 is exactly [18, Proposition 3.17] (with parameter $T$). We therefore keep both target principles $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ explicit in the iteration.
Ryan–Smith also considers the single-parameter instance $\mathsf{PP}(T)$ (Definition 4.1). In general, small-choice hypotheses together with $\mathsf{PP}(T)$ do not suffice to recover global $\mathsf{PP}$: Cohen’s first model satisfies $\mathsf{SVC}^{+}(\mathbb{R})\wedge\mathsf{PP}(\mathbb{R})\wedge\neg\mathsf{PP}$ [18, Proposition 3.22]. Moreover, even adding $\mathsf{AC}_{\mathsf{WO}}$ does not allow one to replace a localized requirement of the form $\mathsf{PP}\!\restriction T$ by a single-parameter assertion $\mathsf{PP}(T)$: the Feferman-style model $N_{\aleph_{1}}$ of Truss satisfies $\mathsf{AC}_{\mathsf{WO}}\wedge\mathsf{SVC}^{+}(\mathcal{P}(\mathbb{R}))\wedge\mathsf{PP}(\mathcal{P}(\mathbb{R}))\wedge\neg\mathsf{PP}$ [18, Proposition 3.23].
Finally, whether $\mathsf{AC}_{\mathsf{WO}}$ follows from $\mathsf{SVC}^{+}(T)\wedge\mathsf{PP}\!\restriction T$ is open in general; see [18, Question 3.18]. This is why $\mathsf{AC}_{\mathsf{WO}}$ remains an explicit forcing target below.
@@ 4.1 The localized $\mathsf{PP}$-package forcing $\mathbb{Q}_{f}$
Fix a transitive model $M\models\mathsf{ZF}+\mathsf{DC}$ extending the Cohen seed $\mathcal{N}$ (in practice, $M$ will be an intermediate stage of our symmetric iteration). Let $T$ be the fixed set defined in $\mathcal{N}$ by Definition 3.27. (Thus $T\in M$, but $T$ need not equal $\mathcal{P}(S)$ as computed in $M$.)
@@
(see [18, Prop. 3.2]). This is the equivalence we invoke in Proposition 5.72.
(For discussion of the hypotheses in Ryan–Smith’s localization (including limitations on weakening them and the redundancy question for $\mathsf{AC}_{\mathsf{WO}}$), see [18, Prop. 3.17, Question 3.18, and Props. 3.22–3.23].)
@@
$\mathbb{Q}_{f}$ is countably closed (i.e. $\omega$-closed) in $M$.
@@
The countable closure of $\mathbb{Q}_{f}$ is proved inside $M$ using $\mathsf{DC}$ (equivalently $\mathsf{AC}_{\omega}$), via “countable union of countable sets is countable”. In a bare $\mathsf{ZF}$ context this need not hold, so Lemma 4.13 should be read as relative to our standing hypothesis $M\models\mathsf{ZF}+\mathsf{DC}$.
@@
Let $M\models\mathsf{ZF}+\mathsf{DC}$ be transitive and let $\mathbb{U}\in M$ be countably closed in $M$ (i.e. every decreasing $\omega$-sequence in $M$ has a lower bound in $M$). Then:
@@
    $\mathbb{U}$ adds no new $\omega$-sequences of ordinals.
@@
    Forcing with $\mathbb{U}$ over $M$ preserves $\mathsf{DC}$.
@@
(1) No new $\omega$-sequences of ordinals. Let $u_{0}\in\mathbb{U}$ and $\dot{a}\in M$ with $u_{0}\Vdash\dot{a}:\check{\omega}\to\mathrm{Ord}$. For each $n<\omega$, let
$D_{n}:=\Bigl\{u\leq u_{0}:\exists\alpha\in\mathrm{Ord}\ (u\Vdash\dot{a}(\check{n})=\check{\alpha})\Bigr\}.$
Each $D_{n}$ is dense below $u_{0}$.
@@
$\Sigma:=\Bigl\{(n,u,\vec{\alpha}):n<\omega,\ u\in\mathbb{U},\ u\leq u_{0},\ \vec{\alpha}\in\mathrm{Ord}^{n},\ u\Vdash\forall k<n\ \dot{a}(\check{k})=\check{\vec{\alpha}(k)}\Bigr\}.$
@@
$(n,u,\vec{\alpha})\ R\ (n+1,v,\vec{\beta})$
iff $v\leq u$, $\vec{\beta}\in\mathrm{Ord}^{n+1}$ extends $\vec{\alpha}$, and $v\Vdash\dot{a}(\check{n})=\check{\vec{\beta}(n)}$. Then $R$ is total on $\Sigma$ by density of $D_{n}$.
By $\mathsf{DC}$ in $M$, there is an $R$-chain $\langle(n,u_{n},\vec{\alpha}_{n}):n<\omega\rangle\in M$ starting from some $(0,u_{0},\emptyset)\in\Sigma$ (note $\mathrm{Ord}^{0}=\{\emptyset\}$). By countable closure, there is $u_{\omega}\in\mathbb{U}$ with $u_{\omega}\leq u_{n}$ for all $n$.
Let $g:=\bigcup_{n<\omega}\vec{\alpha}_{n}$. Since the chain is an element of $M$, Replacement in $M$ yields $g\in M$ and $g:\omega\to\mathrm{Ord}$. By construction, $u_{\omega}\Vdash\forall n\ \dot{a}(\check{n})=\check{g(n)}$, hence $u_{\omega}\Vdash\dot{a}=\check{g}$. Thus $\mathbb{U}$ adds no new $\omega$-sequences of ordinals.
(2) Preservation of $\mathsf{DC}$. Let $u_{0}\in\mathbb{U}$ and $\dot{X},\dot{R},\dot{x}_{0}\in M$ with
$u_{0}\Vdash\ \dot{x}_{0}\in\dot{X}\ \wedge\ (\forall x\in\dot{X})(\exists y\in\dot{X})\ x\,\dot{R}\,y.$
@@
$\Gamma:=\Bigl\{(n,u,\dot{x}):n<\omega,\ u\in\mathbb{U},\ u\leq u_{0},\ \dot{x}\in M,\ u\Vdash\dot{x}\in\dot{X}\Bigr\}.$
@@
$(n,u,\dot{x})\ S\ (n+1,v,\dot{y})$
iff $v\leq u$ and $v\Vdash\dot{y}\in\dot{X}\ \wedge\ \dot{x}\,\dot{R}\,\dot{y}$. Totality of $S$ follows from the premise forced by $u_{0}$.
By $\mathsf{DC}$ in $M$, pick an $S$-chain $\langle(n,u_{n},\dot{x}_{n}):n<\omega\rangle\in M$ starting with $(0,u_{0},\dot{x}_{0})\in\Gamma$. By countable closure, let $u_{\omega}\leq u_{n}$ for all $n$.
@@
$\dot{f}:=\bigl\{\langle\langle\check{n},\dot{x}_{n}\rangle,\mathbbm{1}_{\mathbb{U}}\rangle:n<\omega\bigr\}.$
Since the chain is in $M$, Replacement gives $\dot{f}\in M$. Then $u_{\omega}\Vdash\dot{f}:\check{\omega}\to\dot{X}$ and $u_{\omega}\Vdash\forall n\ \dot{f}(\check{n})\,\dot{R}\,\dot{f}(\check{n+1})$, which is exactly $\mathsf{DC}$ in the extension. ∎
@@
Apply Lemma 4.16 with $\mathbb{U}=\mathbb{Q}_{f}$. Countable closure holds by Lemma 4.13. ∎